Here's another one I need help with:
Every day Andrew wakes up and gets dressed. The probability that he will wear a long-sleeved shirt is 7/10. The probability that he will wear a long-sleeved shirt and shorts is 1/10. The probability that he will wear a long-sleeved shirt OR shorts is 9/10. What is the probability that he will wear shorts?
I think it would be 9/10 less 7/10 plus 1/10 equals 3/10, but that doesn't make logical sense. I think the ANSWER should be 1/10, but if that IS right, how would I arrive at it?

Answer by stanbon(57361)

Every day Andrew wakes up and gets dressed. The probability that he will wear a long-sleeved shirt is 7/10. The probability that he will wear a long-sleeved shirt and shorts is 1/10. The probability that he will wear a long-sleeved shirt OR shorts is 9/10. What is the probability that he will wear shorts?
----
Formula: P(A or B) = P(A) + P(B) - P(A AND B)
----------------------------
P(L OR s) = P(L) + P(s) - P(L AND s)
9/10 = 7/10 + P(s) - 1/10
---
9/10 = 8/10 + P(s)
---
P(shorts) = 1/10
============================
Cheers,
Stan H.
============================